The owner of a music store ordered merchandise with a total value of $11000 from her supplier. She
ordered CDs costing $10 each and DVDs costing $20 each. If she ordered 200 more CDs than DVDs,
how many of each did she order?

Answer by ikleyn(52785)

.

x CDs  and (x-200) DVDs.


Total money equation

    10x + 20*(x-200) = 11000  dollars, total.


Simplify and find x

    10x + 20x - 4000 = 11000

       30x           = 11000 + 4000

       30x           =    15000

         x           =    15000/30 = 500.


ANSWER.  500 CDs  and 500-200 = 300 DVDs.

Solved.